Water flows over hulugan falls, which is 70m high, at an average rate of 2.3 x 10^5 kg/s. If half the po­tential energy of this water were converted into electric en­ergy, how much power would be produced by these falls


1
Expert's answer
2021-05-31T15:38:03-0400

"P_e=W/t=\frac{1}{2}mgh/t=\frac{1}{2}\cdot m_0t\cdot gh/t=\frac{1}{2}\cdot m_0\cdot gh="


"=0.5\cdot2.3\cdot10^5\cdot9.8\cdot70=78.89\cdot10^6 \ (W)=78.89\ (MW)"